OpenStudy (anonymous):

sorry i mean 5/x-33/(x+6)+2=0 sorry

OpenStudy (anonymous):

So it looks like this?: __5____ / (x+6)+2 = 0 x - 33 like tht? (:

OpenStudy (anonymous):

sorry it is 5 divided by x minus 33 divided by (x+6)

OpenStudy (anonymous):

\[\frac{5}{x}-\frac{33}{x+6}+2=0\]